OpenStudy (anonymous):

@fathomboy The diameter of a circle is 42 centimeters. What is the radius of the circle? A. 21 centimeters B. 42 centimeters C. 84 centimeters D. 126 centimeters

OpenStudy (solomonzelman):

The radius is a half of the diameter

OpenStudy (cwrw238):

radius = half of the diameter

OpenStudy (anonymous):

radius = diameter / 2 diameter = 42 radius = diameter / 2 = 21 A

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

What is the solution to the equation? 4(x – 3) = 7x + 24 A. x = 4 B. x = 12 C. x = –12 D. x = –36

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

What is the solution to the inequality? 4x + 7 > –13 A. x < –5 B. x > –5 C. x < 5 D. x > 5

OpenStudy (solomonzelman):

1) subtract 7 from both sides 2) divide both sides by 4 show the work here please.

OpenStudy (anonymous):

I don't get it -_-

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

Which property is shown? (–9 + 17) + 23 = –9 + (17 + 23) A. associative property B. commutative property C. distributive property D. opposite of a sum property

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Which choice lists all the integer solutions of the inequality |x| < 4? A. –4, –3, –2, –1, 0, 1, 2, 3, 4 B. –3, –2, –1, 0, 1, 2, 3 C. –3, –2, –1 D. 0, 1, 2, 3